# Distributed Optimization of Clique-Wise Coupled Problems via Three-Operator Splitting

### v2

@@ 7 Proof of Theorem 2
and thus we get $\xi^{k+1}:=2y^{k+1/2}-y^{k-1/2}-\alpha M^{-1}\nabla_{y}f(y^{k+1/2})+\alpha M^{-1}\nabla_{y}f(y^{k-1/2})\in(I+\alpha M^{-1}\partial g)(y^{k+1})$ by [4, Prop. 16.44]. This inclusion allows us to remove the assumption of smoothness or strong convexity for $g$ or $h$ in [13, 18, 37, 12] because one can evaluate $y^{k+1}$ only with $f$.
We derive the lower side of the inequality in (44) as follows. By the smoothness and strong convexity of $f$, for $\|z^{k}-z^{*}\|^{2}$, [13, Proposition D.4] provides
@@
Next, utilizing a similar calculation to the proof of [13, Proposition D.4] (see the proof of the second upper bound) and the inclusion $\xi^{k+1}\in(I+\alpha M^{-1}\partial g)(y^{k+1})$, we derive the upper-bound of $\|z^{k}-z^{*}\|^{2}+\nu\|\zeta(y^{k-1/2})-\zeta(y^{*})\|^{2}$ as follows. Here, we set $\epsilon=c_{0}/C$, and let $u_{A}^{k+1}\in M^{-1}\partial g(y^{k+1})$ and $u_{A}^{*}\in M^{-1}\partial g(y^{*})$ satisfy $y^{k+1}+\alpha u_{A}^{k*1}=\xi^{k+1}$ and $y^{*}+\alpha u_{A}^{*}=\xi^{*}$, respectively:
@@ Appendix C Proof of Proposition 1
(c) By the chain rule, we have $\frac{\partial}{\partial\mathbf{y}}\hat{f}_{i}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})=\mathbf{D}(\mathbf{D}^{\top}\mathbf{D})^{-1}E_{i}^{\top}\nabla_{x_{i}}\hat{f}_{i}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})$. which gives ((c)).
@@ F.1 Clique-based Projected Gradient Descent (CPGD)
    The fixed points set of $T$ satisfies $\mathrm{Fix}(T)=\mathcal{D}$
@@ F.3 Proof of Proposition 6
a) From Jensen’s inequality and the quasinonexpansiveness of convex projection operators [4], the following inequality holds for any $\mathbf{x},\,\mathbf{w}\in\mathbb{R}^{d}$:
@@
b) $\mathcal{D}\subset\mathrm{Fix}(T)$ holds because $x_{\mathcal{C}_{l}}=P_{\mathcal{D}_{l}}^{Q_{l}}(x_{\mathcal{C}_{l}})$ holds for any $\mathbf{x}\in\mathcal{D}$ and all $l\in\mathcal{Q}_{\mathcal{G}}$. In the following, we prove the converse inclusion $\mathrm{Fix}(T)\subset\mathcal{D}$. Let $\mathbf{w}\in\mathcal{D}$. Then, it suffices to show $\hat{\mathbf{w}}\in\mathrm{Fix}(T)\setminus\{\mathbf{w}\}\Rightarrow\hat{\mathbf{w}}\in\mathcal{D}$. From $\hat{\mathbf{w}}\in\mathrm{Fix}(T)$, we obtain $\hat{w}_{i}=T_{i}(\hat{w}_{\mathcal{N}_{i}})$ for all $i\in\mathcal{N}$. In addition, from Jensen’s inequality and the quasinonexpansiveness of convex projection operators [4], we have
@@
c) For a non-empty closed convex set $\mathcal{D}$ in (85) and $\mathbf{x}\in\mathbb{R}^{d}\setminus\mathcal{D}$, there exists $\hat{l}\in\mathcal{Q}_{\mathcal{G}}$ such that $\|x_{\mathcal{C}_{\hat{l}}}-P_{\mathcal{D}_{\hat{l}}}(x_{\mathcal{C}_{\hat{l}}})\|_{Q_{\hat{l}}}>0$. Hence, for $\hat{l}\in\mathcal{Q}_{\mathcal{G}}$, $\mathbf{x}\in\mathbb{R}^{d}\setminus\mathcal{D}$, and $\mathbf{w}\in\mathcal{D}$, we have $\|x_{\mathcal{C}_{\hat{l}}}-w_{\mathcal{C}_{\hat{l}}}\|^{2}_{Q_{\hat{l}}}>\|P_{\mathcal{D}_{\hat{l}}}^{Q_{\hat{l}}}(x_{\mathcal{C}_{\hat{l}}})-w_{\mathcal{C}_{\hat{l}}}\|^{2}_{Q_{\hat{l}}}$ because
@@
where the last line follows from the projection theorem (see Theorem 3.16 in [4]). Thus, by Jensen’s inequality and the nonexpansiveness of $P_{\mathcal{D}_{l}}^{Q_{l}}$ [4], for any $\mathbf{x}\in\mathbb{R}^{d}\setminus\mathcal{D}$ and $\mathbf{w}\in\mathcal{D}$, we obtain $\|\mathbf{x}-\mathbf{w}\|^{2}=\sum_{l\in\mathcal{Q}_{\mathcal{G}}}\|x_{\mathcal{C}_{l}}-w_{\mathcal{C}_{l}}\|^{2}_{Q_{l}}>\sum_{l\in\mathcal{Q}_{\mathcal{G}}}\|P_{\mathcal{D}_{l}}^{Q_{l}}(x_{\mathcal{C}_{l}})-w_{\mathcal{C}_{l}}\|^{2}_{Q_{l}}\geq\sum_{i=1}^{n}\|\sum_{l\in\mathcal{Q}_{\mathcal{G}}^{i}}\frac{1}{|\mathcal{Q}_{\mathcal{G}}^{i}|}E_{l,i}P_{\mathcal{D}_{l}}^{Q_{l}}(x_{\mathcal{C}_{l}})=\|T(\mathbf{x})-\mathbf{w}\|^{2}.$ Hence, $\|T(\mathbf{x})-\mathbf{w}\|<\|\mathbf{x}-\mathbf{w}\|$ for any $\mathbf{x}\in\mathbb{R}^{d}\setminus\mathcal{D}$ and $\mathbf{w}\in\mathcal{D}$.
d) For $\mathbf{x}\in\mathbb{R}^{d}$, we define $\{a_{k}\}$ as $a_{k+1}=T(a_{k})$ with $a_{0}=\mathbf{x}$. Then, we obtain $\lim_{k\to\infty}a_{k+1}=\lim_{k\to\infty}T(a_{k})$. Thus, from the continuity of $T$ shown in Proposition 6a, we have $T^{\infty}(x)=\lim_{k\to\infty}a_{k+1}=T(\lim_{k\to\infty}a_{k})=T(T^{\infty}(x))$. Hence, Proposition 6b yields $T^{\infty}(x)\in\mathrm{Fix}(T)=\mathcal{D}$.
@@ F.4 Proof of Theorems 3c and 4
$=$ $G_{\mathbf{w}}(\boldsymbol{\xi})+\frac{1}{\alpha}(\mathbf{x}-T(\mathbf{w}))^{\top}(\mathbf{x}-\boldsymbol{\xi})-\frac{1}{2\alpha}\|\mathbf{x}-T(\mathbf{w})\|$

### v3

@@ 7 Proof of Theorem 2
and thus we get $\xi^{k+1}:=2y^{k+1/2}-y^{k-1/2}-\alpha M^{-1}\nabla_{y}f(y^{k+1/2})+\alpha M^{-1}\nabla_{y}f(y^{k-1/2})\in(I+\alpha M^{-1}\partial g)(y^{k+1})$ by [30, Prop. 16.44]. This inclusion allows us to remove the assumption of smoothness or strong convexity for $g$ or $h$ in [12, 20, 21, 22] because one can evaluate $y^{k+1}$ only with $f$, without $z^{k}$.
We derive the lower side of the inequality in (43) as follows. By the smoothness and strong convexity of $f$, for $\|z^{k}-z^{*}\|^{2}$, [12, Proposition D.4] provides
@@
Next, utilizing a similar calculation to the proof of [12, Proposition D.4] (see the proof of the second upper bound) and the inclusion $\xi^{k+1}\in(I+\alpha M^{-1}\partial g)(y^{k+1})$, we derive the upper-bound of $\|z^{k}-z^{*}\|^{2}+\nu\|\zeta(y^{k-1/2})-\zeta(y^{*})\|^{2}$ as follows. Here, we set $\epsilon=c_{0}/C$, and let $u_{A}^{k+1}\in M^{-1}\partial g(y^{k+1})$ and $u_{A}^{*}\in M^{-1}\partial g(y^{*})$ satisfy $y^{k+1}+\alpha u_{A}^{k*1}=\xi^{k+1}$ and $y^{*}+\alpha u_{A}^{*}=\xi^{*}$, respectively:
@@ Appendix C Proof of Proposition 1
(c) By the chain rule, we have $\frac{\partial}{\partial\mathbf{y}}\hat{f}_{i}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})=\mathbf{D}(\mathbf{D}^{\top}\mathbf{D})^{-1}E_{i}^{\top}\nabla_{x_{i}}\hat{f}_{i}(E_{i}(\mathbf{D}^{\top}\mathbf{D})^{-1}\mathbf{D}^{\top}\mathbf{y})$. which gives (17).
@@ F.1 Clique-based Projected Gradient Descent (CPGD)
    The fixed points set of $T$ satisfies $\mathrm{Fix}(T)=\mathcal{D}$.
@@ F.3 Proof of Proposition 6
(a) From Jensen’s inequality and the quasinonexpansiveness of convex projection operators [30], the following inequality holds for any $\mathbf{x},\,\mathbf{w}\in\mathbb{R}^{d}$:
@@
(b) $\mathcal{D}\subset\mathrm{Fix}(T)$ holds because $x_{\mathcal{C}_{l}}=P_{\mathcal{D}_{l}}^{Q_{l}}(x_{\mathcal{C}_{l}})$ holds for any $\mathbf{x}\in\mathcal{D}$ and all $l\in\mathcal{Q}_{\mathcal{G}}$. In the following, we prove the converse inclusion $\mathrm{Fix}(T)\subset\mathcal{D}$. Let $\mathbf{w}\in\mathcal{D}$. Then, it suffices to show $\hat{\mathbf{w}}\in\mathrm{Fix}(T)\setminus\{\mathbf{w}\}\Rightarrow\hat{\mathbf{w}}\in\mathcal{D}$. From $\hat{\mathbf{w}}\in\mathrm{Fix}(T)$, we obtain $\hat{w}_{i}=T_{i}(\hat{w}_{\mathcal{N}_{i}})$ for all $i\in\mathcal{N}$. In addition, from Jensen’s inequality and the quasinonexpansiveness of convex projection operators [30], we have
@@
(c) For a non-empty closed convex set $\mathcal{D}$ in (84) and $\mathbf{x}\in\mathbb{R}^{d}\setminus\mathcal{D}$, there exists $\hat{l}\in\mathcal{Q}_{\mathcal{G}}$ such that $\|x_{\mathcal{C}_{\hat{l}}}-P_{\mathcal{D}_{\hat{l}}}(x_{\mathcal{C}_{\hat{l}}})\|_{Q_{\hat{l}}}>0$. Hence, for $\hat{l}\in\mathcal{Q}_{\mathcal{G}}$, $\mathbf{x}\in\mathbb{R}^{d}\setminus\mathcal{D}$, and $\mathbf{w}\in\mathcal{D}$, we have $\|x_{\mathcal{C}_{\hat{l}}}-w_{\mathcal{C}_{\hat{l}}}\|^{2}_{Q_{\hat{l}}}>\|P_{\mathcal{D}_{\hat{l}}}^{Q_{\hat{l}}}(x_{\mathcal{C}_{\hat{l}}})-w_{\mathcal{C}_{\hat{l}}}\|^{2}_{Q_{\hat{l}}}$ because
@@
where the last line follows from the projection theorem (see Theorem 3.16 in [30]). Thus, by Jensen’s inequality and the nonexpansiveness of $P_{\mathcal{D}_{l}}^{Q_{l}}$ [30], for any $\mathbf{x}\in\mathbb{R}^{d}\setminus\mathcal{D}$ and $\mathbf{w}\in\mathcal{D}$, we obtain $\|\mathbf{x}-\mathbf{w}\|^{2}=\sum_{l\in\mathcal{Q}_{\mathcal{G}}}\|x_{\mathcal{C}_{l}}-w_{\mathcal{C}_{l}}\|^{2}_{Q_{l}}>\sum_{l\in\mathcal{Q}_{\mathcal{G}}}\|P_{\mathcal{D}_{l}}^{Q_{l}}(x_{\mathcal{C}_{l}})-w_{\mathcal{C}_{l}}\|^{2}_{Q_{l}}\geq\sum_{i=1}^{n}\|\sum_{l\in\mathcal{Q}_{\mathcal{G}}^{i}}\frac{1}{|\mathcal{Q}_{\mathcal{G}}^{i}|}E_{l,i}P_{\mathcal{D}_{l}}^{Q_{l}}(x_{\mathcal{C}_{l}})\|^{2}=\|T(\mathbf{x})-\mathbf{w}\|^{2}.$ Hence, $\|T(\mathbf{x})-\mathbf{w}\|<\|\mathbf{x}-\mathbf{w}\|$ for any $\mathbf{x}\in\mathbb{R}^{d}\setminus\mathcal{D}$ and $\mathbf{w}\in\mathcal{D}$.
(d) For $\mathbf{x}\in\mathbb{R}^{d}$, we define $\{a_{k}\}$ as $a_{k+1}=T(a_{k})$ with $a_{0}=\mathbf{x}$. Then, we obtain $\lim_{k\to\infty}a_{k+1}=\lim_{k\to\infty}T(a_{k})$. Thus, from the continuity of $T$ shown in Proposition 6a, we have $T^{\infty}(x)=\lim_{k\to\infty}a_{k+1}=T(\lim_{k\to\infty}a_{k})=T(T^{\infty}(x))$. Hence, Proposition 6b yields $T^{\infty}(x)\in\mathrm{Fix}(T)=\mathcal{D}$.
@@ F.4 Proof of Theorems 3c and 4
$=$ $G_{\mathbf{w}}(\boldsymbol{\xi})+\frac{1}{\alpha}(\mathbf{x}-T(\mathbf{w}))^{\top}(\mathbf{x}-\boldsymbol{\xi})-\frac{1}{2\alpha}\|\mathbf{x}-T(\mathbf{w})\|^{2}$